What does the x represent in this situation 3x=-4y-21 and 5x+7y=39?
Answer by stanbon(75887)

What does the x represent in this situation
3x = -4y-21
5x+7y =39
----
Rearrange:
3x + 4y = -21
5x + 7y = 39
-----
Multiply thru top by 7 and bottom by 4:
21x + 28y = -147
20x + 28y = 156
---
Subtract and solve for "x":
x = -303
=========
Cheers,
Stan H.
=========